would you please solve following question?
The population of a town increased by 50/3%
in 1990 and by 10% in 1991. If the population of the town at the beginning of 1990 was 30,000, what was it at the end of 1991?
-Thank you.

Answer by Theo(13342)

at the beginning of 1990, the population was 30000.

it grew by 50/3% in that year.

50/3% = 50/3/100 = 50/300 = 5/30 = 1/6.

1/6 * 30000 = 5000

at the beginning of 1991, the population was 30000 + 5000 = 35000.

it grew by 10% in that year.

10% = 10/100 = 1/10.

1/10 * 35000 = 3500.

at the beginning of 1992, the population was 35000 + 3500 = 38500.

the beginning of 1991 is the same as the end of 1990.
the beginning of 1992 is the same as the end of 1991.